OpenStudy (anonymous):

There are 6 red marbles, 9 blue marbles, and 10 green marbles in a bag. What is the theoretical probability of randomly drawing a red marble and then a green marble?

OpenStudy (anonymous):

depends entirely on whether you replace the marble or not after the first one is drawn

OpenStudy (anonymous):

The first one is not replaced.

OpenStudy (anonymous):

oh wait NOT replaced lets go slow

OpenStudy (anonymous):

what is the probability the first one is red?

OpenStudy (anonymous):

i hope that is more or less obvious it is the ratio of the number or red marbles to the total number of marbles

OpenStudy (anonymous):

24%?

OpenStudy (anonymous):

There are 4 answers: 10% 9.6% 16% 64%

OpenStudy (anonymous):

not really interested in percents before we compute these numbers we need fractions first, then we can turn the final answer in to a percent

OpenStudy (anonymous):

\[\frac{ 24 }{100}?\]

OpenStudy (anonymous):

lets go slow how many red marbles are there?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes, and how many marbles are there total?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

right so the probability of picking a red marble first is \[\frac{6}{25}\]

OpenStudy (anonymous):

now the red marble is chosen, and apparently not replaced how many green marbles are there ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

and how many marbles are left in the bag?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

so the probability of picking a green marble once one red one is removes is?

OpenStudy (anonymous):

\[\frac{ 10 }{ 24 }\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to compute the probability that both things occur, multiply

OpenStudy (anonymous):

\[\frac{6}{25}\times \frac{10}{24}\]

OpenStudy (anonymous):

\[\frac{ 60 }{ 600 }\]

OpenStudy (anonymous):

better known on planet earth as \(\frac{1}{10}\)

OpenStudy (anonymous):

or \(0.1\) or even \(10\%\)

OpenStudy (anonymous):

So that's the answer?